# AIDOVECL: AI-generated Dataset of Outpainted Vehicles for Eye-level Classification and Localization

### v2

@@ Abstract
Image labeling is a critical bottleneck in the development of computer vision technologies, often constraining the potential of machine learning models due to the time-intensive nature of manual annotations. This work introduces a novel approach that leverages outpainting to mitigate the problem of annotated data scarcity by generating artificial contexts and annotations, significantly reducing manual labeling efforts. We apply this technique to a particularly acute challenge in autonomous driving, urban planning, and environmental monitoring: the lack of diverse, eye-level vehicle images in desired classes. Our dataset comprises AI-generated vehicle images obtained by detecting and cropping vehicles from manually selected seed images, which are then outpainted onto larger canvases to simulate varied real-world conditions. The outpainted images include detailed annotations, providing high-quality ground truth data. Advanced outpainting techniques and image quality assessments ensure visual fidelity and contextual relevance. Ablation results show that incorporating AIDOVECL improves overall detection performance by up to 10%, and delivers gains of up to 40% in settings with greater diversity of context, object scale, and placement, with underrepresented classes achieving up to 50% higher true positives. AIDOVECL enhances vehicle detection by augmenting real training data and supporting evaluation across diverse scenarios. By demonstrating outpainting as an automatic annotation paradigm, it offers a practical and versatile solution for building fine-grained datasets with reduced labeling effort across multiple machine learning domains. The code and links to datasets used in this study are available for further research and replication at https://github.com/amir-kazemi/aidovecl.
@@ 1 Introduction
In recent years, the field of computer vision has undergone significant expansion, leading to transformative changes across various domains such as autonomous driving, urban planning, and environmental monitoring. These advanced technologies are set to revolutionize transportation by reducing traffic accidents, easing congestion, and mitigating urban air pollution through improved vehicle operations (dilek2023computer; fan2024enhancing; mavrokapnidis2021community). Central to these advancements is the robustness and accuracy of vehicle classification and localization algorithms, which depend critically on the quality and diversity of the training datasets available. These technologies are continuously being developed to enhance vehicle detection for various applications, ranging from the detection of polluting vehicles (nisha2025improved) to the handling of corner cases in autonomous driving (lu2024realistic). Such advancements underscore the critical need for a broad spectrum of vehicle images within specific categories to support these technologies.
Vision-based localization methods, despite their reliance on optimal lighting conditions and susceptibility to obstructions, stand out due to their robust performance in familiar settings. These methods excel by recognizing visual features to enhance navigation and obstacle detection, providing rich contextual data that not only enhances system functionality but also ensures a comprehensive environmental understanding (lu2021real). Nonetheless, the efficacy of these classification and localization models hinges on access to varied and extensive datasets. Public datasets frequently lack adequate eye-level vehicle representations—essential for autonomous driving and roadside surveillance applications. Furthermore, these datasets often do not include detailed or desired vehicle categorizations, thereby limiting their practical utility. This is further exacerbated by the presence of smaller or partial vehicles in the background which are not annotated. All these shortcomings can adversely affect the performance of vision-based localization machine learning models tasked with operating in dynamic, real-world settings.
@@
The contributions of this dataset are manifold, significantly enhancing the adaptability and performance of machine learning models. The use of outpainting not only increases the diversity of the dataset but also simulates real-world conditions where vehicles appear at various sizes in various parts of an image. This simulation is crucial for developing robust algorithms capable of accurate vehicle classification and localization under diverse operational scenarios. Each image within the dataset is automatically annotated with detailed bounding box coordinates, providing valuable ground truth data for training and evaluation purposes. Also, seed images can readily be categorized into desired classes manually before populating them by outpainting. We may use AIDOVECL as a method to augment real vehicle image datasets which suffer from data scarcity or class imbalance. Our comprehensive ablation study across multiple detection architectures clearly demonstrates the effectiveness of incorporating AIDOVECL, yielding improvements of up to 10% in overall detection performance. The performance gain can rise up to 40% in settings characterized by greater diversity of context, object scale, and placement, with underrepresented classes achieving up to 50% higher true positives. These results highlight AIDOVECL’s primary value in augmenting real training data to strengthen vehicle detection, while also serving as a reliable supplement for testing in diverse scenarios. Beyond these performance gains, the study also offers insight into how well existing quality assessment measures capture perceptual and semantic aspects of generated images, particularly in prompt-guided image-to-image tasks like outpainting. By combining advanced detection techniques with state-of-the-art generative models and quality assessment tools, our pipeline and dataset facilitates progress in autonomous driving, traffic analysis, and urban planning. This capability is especially useful in scenarios requiring custom fine-grained classes, reducing labeling effort by decoupling manual classification from object localization and context variation.
@@
The rest of the paper is structured as follows: The Background section reviews significant vehicle image datasets, identifying gaps that our approach aims to fill, and also includes an overview of related inpainting and augmentation techniques. The Methods and Setup section describes the process of creating seed vehicle images, generating synthetic data through outpainting, and assessing image quality, along with the training and evaluation of the model for vehicle classification and localization. In Results and Discussion, we present the outcomes of our image generation and quality assessment, illustrated with examples of outpainted vehicle images, and discuss the performance of the model in classification and localization tasks. Finally, the Conclusion summarizes our contributions, highlighting the effectiveness of outpainting as an automatic annotation approach, along with its limitation and possible future works.
@@ 2.1 Public Vehicle Image Datasets
Vehicle datasets are essential for improving model accuracy in vehicle recognition across different environments. Despite their importance, these datasets frequently encounter challenges such as limited context around vehicles, broad classifications that obscure finer distinctions, and suboptimal capture angles. These limitations can severely affect tasks requiring detailed environmental information, precise vehicle identification, or particular viewing angles. We briefly review major vehicle datasets to highlight these shortcomings and identify the need for more customizable datasets.
The Stanford Cars dataset contains over 16,000 images of 196 classes of cars (krause20133d), covering a wide range of makes, models, and manufacturing years. This dataset is particularly valued for its high-quality images and annotations, making it a standard benchmark in vehicle recognition and classification tasks. However, many images include unlabeled vehicles in the background, and in some instances, the images do not provide enough context, being either too close to the bounding box or having vehicles surrounded by blank spaces, which is not ideal for training highly accurate localization models. Besides, the dataset does not include heavy trucks and buses.
The Miovision Traffic Camera Dataset (MIO-TCD) dataset (luo2018mio) is another significant resource utilized in vehicle recognition and traffic scene understanding. It comprises a diverse set of traffic-related images designed to facilitate the training and evaluation of various algorithms on mobile platforms. While the dataset offers a broad variety of vehicle types (10 vehicle classes) and traffic scenarios, its major drawback is the lack of near eye-level perspectives. Most images are captured from elevated angles, which can limit the dataset’s applicability for applications requiring a pedestrian’s, driver’s, or roadside point of view.
The Common Objects in Context (COCO) dataset (lin2014microsoft) is a cornerstone in computer vision, known for its robust features in object detection, segmentation, and captioning across diverse scenes. It excels due to extensive annotations, including precise segmentation masks essential for object localization. However, its vehicle categorization is limited to just three classes: car, bus, and truck. This broad classification can be inadequate for specialized applications needing detailed vehicle recognition, such as advanced safety and surveillance systems.
@@ 2.2 Inpainting Methods
Image inpainting is a computational technique used to restore and manipulate images, primarily for removing unwanted objects or artifacts. It has applications across various domains, including security, where it plays a critical role in ensuring the integrity of visual information by eliminating potentially compromising elements from shared images. Originally employed for repairing old or damaged images, it has evolved to address a range of distortions such as text, noise, scratches, and lines. While serving as a valuable tool for enhancing image fidelity, it also presents challenges in forgery detection due to the potential for sophisticated manipulation techniques to conceal traces of editing (elharrouss2020image).
@@
A variety of strategies are employed to reconstruct missing or corrupted image areas, ranging from progressive (8578675; sagong2019pepsi; guo2021jpgnet; zhang2018semantic; li2020recurrent; yi2020contextual) and structural information-guided (nazeri2019edgeconnect; yu2019free; liao2020guidance; han2019deep; yang2020learning) to attention-based (zeng2020high; zhao2020uctgan; liu2019coherent; wan2021high; yu2021diverse) approaches. These strategies leverage iterative filling processes, structural and spatial data, and controlled information propagation via masks to mitigate color inconsistencies. Pluralistic approaches address the inherent complexity of inpainting by generating diverse possible reconstructions for each image segment (xiang2023deep). Extending these foundational strategies, various deep learning methods utilize frameworks such as Generative Adversarial Networks (GAN) (goodfellow2020generative) and Variational Autoencoder (VAE) (kingma2013auto) integrated with Convolutional Neural Networks (CNN) (lecun1995convolutional) or Recurrent Neural Networks (RNN) (graves2013generating) to advance inpainting for a diverse set of applications (zhang2023image).
@@
The recent introduction of Latent Diffusion Models (LDM) represents a significant advancement for the inpainting task (rombach2022high). This method leverages latent diffusion processes to synthesize high-resolution images, focusing on intricate details and structural coherence that outpace traditional models. Unlike previous techniques that often concentrated on replicating textures and patterns, LDMs excel in generating and processing contextual and structural details crucial for complex inpainting tasks. The diffusion process iteratively refines images through transitions across latent representations, enabling effective handling of various distortions without extensive manual parameter tuning. This breakthrough offers a powerful, scalable solution for high-fidelity image inpainting, setting a new benchmark for inpainting tools and markedly enhancing both the quality and realism of inpainted areas in academic and practical applications. Outpainting generates content beyond an image’s original boundaries using similar pixel reconstruction techniques as inpainting. Consequently, we employ LDM (rombach2022high), an advanced method with established pretrained models, to leverage these capabilities.
@@ 2.3 Related Data Augmentation Methods
Data augmentation is a common technique in machine learning, used to enhance the size and diversity of training datasets, thereby improving the generalizability of models. Traditional methods like flipping and rotation have long been staples, but more sophisticated techniques are necessary to tackle the complex demands of modern AI systems. While augmentation methods such as mosaic and mixup have proved to be promising, the advent of outpainting using diffusion models for object augmentation in diverse environments remains fairly unexplored.
@@
Mosaic data augmentation, introduced in the YOLOv4 paper (bochkovskiy2020yolov4), represents a significant advancement in training deep learning models for object detection. This technique involves combining four or more different training images into a single composite image. By doing so, it exposes the model to a more diverse array of object scales, positions, and contexts within a single training example. This approach not only helps in enhancing the detection capabilities of the model, especially for smaller objects, but also significantly boosts its ability to generalize to various real-world scenarios. The inclusion of multiple scenes in one image helps in simulating a more complex and varied visual environment, which is crucial for improving the robustness and accuracy of object detection systems. This method has proven to be especially effective in handling datasets where objects appear in dense configurations and diverse backgrounds.
@@
Mixup data augmentation (zhang2018mixup), significantly enhances the training of neural networks for image classification and related tasks. This method operates by blending two images and their corresponding labels, creating a new, synthetic training example. Mixup achieves this blending by a convex mixture of two images and their corresponding labels, with weights that are randomly determined yet constrained to sum to one. By doing this, mixup encourages the model to adopt simpler, linear behaviors between training examples, mitigating issues of overfitting and improving model generalization. This linear interpolation between examples provides smoother estimates of uncertainty, which is particularly beneficial in scenarios where models must handle novel or ambiguous inputs. The inherent noise introduced through this process also helps to robustify the model against small perturbations in input space, thus making it more adaptable and robust in practical applications.
@@
While Mosaic and Mixup augmentation techniques have made significant strides in enhancing neural network robustness, they also present notable limitations. Mosaic can sometimes result in overly complex images that confuse the model rather than help it learn, especially when objects from different images do not naturally coexist. Similarly, mixup might lead to ambiguous labels if the images combined are too distinct, potentially misleading the model during training. Furthermore, methods like copy-paste (ghiasi2021simple) and x-paste (zhao2023x), which rely on precise segmentation masks and may not be as useful with just bounding box annotations, often face challenges in integrating objects seamlessly into new scenes, leading to unnatural edges or misplaced objects. In contrast, outpainting using diffusion models emerges as a more promising avenue. These models can generate highly realistic and contextually varied synthetic data by extending the canvas of existing images, creating entirely new scenarios that maintain the integrity and realism of the original content. Additionally, the availability of pretrained prompt-guided diffusion models allows for efficient deployment, significantly speeding up the development cycle and utilizing resources merely for inference. This method not only addresses the limitations of conventional augmentation techniques but also provides a scalable solution to train more robust and adaptable AI systems.
@@ 3.1 Creating Seed Vehicle Images
The initial step involves gathering a diverse collection of images of vehicles across the desired categories. This process is manual and entails selecting images where vehicles are presented in a near-eye-level perspective, ensuring that each vehicle is well-separated from smaller background vehicles, if present. Additionally, it is crucial that vehicles are fully visible within the frame, as partial representations result in information loss. For this analysis, our collection comprises over 11,000 vehicle images from various sources (including Stanford Cars) (krause20133d; peng2019moment; imagenet-object-localization-challenge; sedan-cars_dataset; bignosethethird2024uktruckbrands; gerritonagoodday2024vehiclebrandscraping; bus_photos_dataset; koul2024vehicledataset) which are utilized exclusively for academic and non-commercial research purposes.
@@
We select the aforementioned detection model by a consensus voting approach, where each model in a predefined ensemble—FCOS (tian2019fcos), RetinaNet (lin2017focal), SSD (liu2016ssd), MaskRCNN (he2017mask), and FasterRCNN (ren2016faster)—detects the largest bounding box for vehicles in a given image. These detections are then compared pairwise using the Intersection over Union (IoU) metric, where each bounding box gains a vote for every other box it significantly overlaps with (as defined by a threshold of 0.95). The model with the highest votes is selected as the primary detector, with the order indicating backup priorities in case the primary model fails. This method leverages multiple models to increase the reliability of the detection by confirming the presence and location of objects through mutual agreement.

### v3

@@ Abstract
Image labeling is a critical bottleneck in the development of computer vision technologies, often constraining machine learning performance due to the time-intensive nature of manual annotations. This work introduces a novel approach that leverages outpainting to mitigate annotated data scarcity by generating artificial contexts and annotations, significantly reducing labeling efforts. We apply this technique to a particularly acute challenge in autonomous driving, urban planning, and environmental monitoring: the lack of diverse, eye-level vehicle images from desired classes. Our dataset comprises AI-generated vehicle images obtained by detecting and cropping vehicles from manually selected seed images, which are then outpainted onto larger canvases to simulate varied real-world conditions. The outpainted images include detailed annotations, providing high-quality ground truth data. Advanced outpainting techniques and image quality assessments ensure visual fidelity and contextual relevance. Ablation results show that incorporating AIDOVECL improves overall detection performance by up to about 10%, and delivers gains of up to about 40% in settings with greater diversity of context, object scale, and placement, with underrepresented classes achieving up to about 50% higher true positives. AIDOVECL enhances vehicle detection by augmenting real training data and supporting evaluation across diverse scenarios. By demonstrating outpainting as an automatic annotation paradigm, it offers a practical and versatile solution for building fine-grained datasets with reduced labeling effort across multiple machine learning domains. The code and links to datasets are available for further research and replication at https://github.com/amir-kazemi/aidovecl.
@@ 1 Introduction
In recent years, the field of computer vision has undergone significant expansion, leading to transformative changes across various domains such as autonomous driving, urban planning, and environmental monitoring. These advanced technologies may catalyze improvements in transportation, leading to reduced traffic accidents, easing congestion, and mitigating urban air pollution through improved vehicle operations (Dilek and Dener, 2023; Fan et al., 2024; Mavrokapnidis et al., 2021). Central to these advancements is the robustness and accuracy of vehicle classification and localization algorithms, which depend critically on the quality and diversity of the training datasets available. These technologies are continuously being developed to enhance vehicle detection for various applications, ranging from the detection of polluting vehicles (Nisha et al., 2025) to the handling of corner cases in autonomous driving (Lu et al., 2024). Such advancements underscore the critical need for a broad spectrum of vehicle images within specific categories to support these technologies.
Vision-based localization methods, despite their reliance on optimal lighting conditions and susceptibility to obstructions, stand out due to their robust performance in familiar settings. These methods excel by recognizing visual features to enhance navigation and obstacle detection, providing rich contextual data that not only enhances system functionality but also ensures a comprehensive environmental understanding (Lu et al., 2021). Nonetheless, the efficacy of these classification and localization models hinges on access to varied and extensive datasets. Public datasets frequently lack adequate eye-level vehicle representations—essential for autonomous driving and roadside surveillance applications. Furthermore, these datasets often do not include detailed or desired vehicle categorizations, thereby limiting their practical utility. This is further exacerbated by the presence of smaller or partial vehicles in the background which are not annotated. All these shortcomings can adversely affect the performance of vision-based localization machine learning models tasked with operating in dynamic, real-world settings.
@@
The contributions of this dataset are manifold, significantly enhancing the adaptability and performance of machine learning models. The use of outpainting not only increases the diversity of the dataset but also simulates real-world conditions where vehicles appear at varying scales and locations within an image. This simulation is crucial for developing robust algorithms capable of accurate vehicle classification and localization under diverse operational scenarios. Each image within the dataset is automatically annotated with detailed bounding box coordinates, providing valuable ground truth data for training and evaluation purposes. Also, seed images can readily be categorized into desired classes manually before populating them by outpainting. We may use AIDOVECL as a method to augment real vehicle image datasets which suffer from data scarcity or class imbalance. Our comprehensive ablation study across multiple detection architectures clearly demonstrates the effectiveness of incorporating AIDOVECL, yielding improvements of up to 10% in overall detection performance. The performance gain can rise up to 40% in settings characterized by greater diversity of context, object scale, and placement, with underrepresented classes achieving up to 50% higher true positives. These results highlight AIDOVECL’s primary value in augmenting real training data to strengthen vehicle detection, while also serving as a reliable supplement for testing in diverse scenarios. Beyond these performance gains, the study also offers insight into how well existing quality assessment measures capture perceptual and semantic aspects of generated images, particularly in prompt-guided image-to-image tasks like outpainting. By combining advanced detection techniques with state-of-the-art generative models and quality assessment tools, our pipeline and dataset facilitate progress in autonomous driving, traffic analysis, and urban planning. This capability is especially useful in scenarios requiring custom fine-grained classes, reducing labeling effort by decoupling manual classification from object localization and context variation.
@@
The rest of the paper is structured as follows: The Background section reviews significant vehicle image datasets, identifying gaps that our approach aims to fill, and also includes an overview of related inpainting and augmentation techniques. The Methods and Setup section describes the process of creating seed vehicle images, generating synthetic data through outpainting, and assessing image quality, along with the training and evaluation of the model for vehicle classification and localization. In Results and Discussion, we present the outcomes of our image generation and quality assessment, illustrated with examples of outpainted vehicle images, and discuss the performance of the model in classification and localization tasks. Finally, the Conclusion summarizes our contributions, highlighting the effectiveness of outpainting as an automatic annotation approach, along with its limitations and possible future work.
@@ 2.1 Public Vehicle Image Datasets
Vehicle datasets are essential for improving model accuracy in vehicle recognition in different environments. Despite their importance, these datasets frequently encounter challenges such as limited context around vehicles, broad classifications that obscure finer distinctions, and suboptimal capture angles. These limitations can severely affect tasks requiring detailed environmental information, precise vehicle identification, or particular viewing angles. We briefly review major vehicle datasets to highlight these shortcomings and identify the need for datasets that are more customizable.
The Stanford Cars dataset contains over 16,000 images of 196 classes of cars (Krause et al., 2013), covering a wide range of makes, models, and manufacturing years. This dataset is particularly valued for its high-quality images and annotations, making it a standard benchmark in vehicle recognition and classification tasks. However, many images include unlabeled vehicles in the background, and in some instances, the images do not provide much context, being either too close to the bounding box or having vehicles surrounded by blank spaces, which is not ideal for training highly accurate localization models. Additionally, the dataset does not include heavy trucks and buses.
The Miovision Traffic Camera Dataset (MIO-TCD) (Luo et al., 2018) is another significant resource utilized in vehicle recognition and traffic scene understanding. It comprises a diverse set of traffic-related images designed to facilitate the training and evaluation of various algorithms on mobile platforms. While the dataset offers a broad variety of vehicle types (10 vehicle classes) and traffic scenarios, its major drawback is the lack of near eye-level perspectives. Most images are captured from elevated angles, which can limit the dataset’s applicability for applications requiring a pedestrian’s, driver’s, or roadside point of view.
The Common Objects in Context (COCO) dataset (Lin et al., 2014) is a cornerstone in computer vision, known for its robust features in object detection, segmentation, and captioning across diverse scenes. It excels due to extensive annotations, including precise segmentation masks essential for object localization. However, its vehicle categorization is limited to just three classes: car, bus, and truck. This broad classification can be inadequate for specialized applications needing detailed vehicle recognition, such as advanced safety and surveillance systems.
@@ 2.2 Inpainting Methods
Image inpainting is a computational technique used to restore and manipulate images, primarily for removing unwanted objects or artifacts. It has applications across various domains, including security, where it plays a critical role in ensuring the integrity of visual information by eliminating potentially compromising elements from shared images. Originally employed for repairing old or damaged images, it has evolved to address a range of distortions such as text, noise, scratches, and lines. While serving as a valuable tool for enhancing image fidelity, it also presents challenges in forgery detection due to the potential for sophisticated manipulation techniques to conceal traces of editing (Elharrouss et al., 2020).
@@
A variety of strategies are employed to reconstruct missing or corrupted image areas, ranging from progressive (Yu et al., 2018; Sagong et al., 2019; Guo et al., 2021; Zhang et al., 2018a; Li et al., 2020; Yi et al., 2020) and structural information-guided (Nazeri et al., 2019; Yu et al., 2019; Liao et al., 2020; Han et al., 2019; Yang et al., 2020) to attention-based (Zeng et al., 2020; Zhao et al., 2020; Liu et al., 2019; Wan et al., 2021; Yu et al., 2021) approaches. These strategies leverage iterative filling processes, structural and spatial data, and controlled information propagation via masks to mitigate color inconsistencies. Pluralistic approaches address the inherent complexity of inpainting by generating diverse possible reconstructions for each image segment (Xiang et al., 2023). Extending these foundational strategies, various deep learning methods utilize frameworks such as Generative Adversarial Networks (GAN) (Goodfellow et al., 2020) and Variational Autoencoder (VAE) (Kingma and Welling, 2013) integrated with Convolutional Neural Networks (CNN) (LeCun et al., 1995) or Recurrent Neural Networks (RNN) (Graves, 2013) to advance inpainting for a diverse set of applications (Zhang et al., 2023b).
@@
The recent introduction of Latent Diffusion Models (LDM) represents a significant advancement for the inpainting task (Rombach et al., 2022). This method leverages latent diffusion processes to synthesize high-resolution images, focusing on intricate details and structural coherence that outpace traditional models. Unlike previous techniques that often concentrated on replicating textures and patterns, LDMs excel in generating and processing contextual and structural details crucial for complex inpainting tasks. The diffusion process iteratively refines images through transitions across latent representations, enabling effective handling of various distortions without extensive manual parameter tuning. This breakthrough offers a powerful, scalable solution for high-fidelity image inpainting, setting a new benchmark for inpainting tools and markedly enhancing both the quality and realism of inpainted areas in academic and practical applications. Outpainting generates content beyond an image’s original boundaries using similar pixel reconstruction techniques as inpainting. Consequently, we employ LDM (Rombach et al., 2022), an advanced method with established pretrained models, to leverage these capabilities.
@@ 2.3 Related Data Augmentation Methods
Data augmentation is a common technique in machine learning, used to enhance the size and diversity of training datasets, thereby improving the generalizability of models. Traditional methods like flipping and rotation have long been staples, but more sophisticated techniques are necessary to tackle the complex demands of modern AI systems. While augmentation methods such as mosaic and mixup have proved to be promising, the advent of outpainting using diffusion models for object augmentation in diverse environments remains relatively unexplored.
@@
Mosaic data augmentation, introduced in the YOLOv4 paper (Bochkovskiy et al., 2020), represents a significant advancement in training deep learning models for object detection. This technique involves combining four or more different training images into a single composite image. By doing so, it exposes the model to a more diverse array of object scales, positions, and contexts within a single training example. This approach not only helps in enhancing the detection capabilities of the model, especially for smaller objects, but also significantly boosts its ability to generalize to various real-world scenarios. The inclusion of multiple scenes in one image helps in simulating a more complex and varied visual environment, which is crucial for improving the robustness and accuracy of object detection systems. This method has proven to be especially effective in handling datasets where objects appear in dense configurations and diverse backgrounds.
@@
Mixup data augmentation (Zhang et al., 2018b), significantly enhances the training of neural networks for image classification and related tasks. This method generates a synthetic training sample by combining two images and their corresponding labels using a convex mixture. By doing this, mixup encourages the model to adopt simpler, linear behaviors between training examples, mitigating issues of overfitting and improving model generalization. This linear interpolation between examples provides smoother estimates of uncertainty, which is particularly beneficial in scenarios where models must handle novel or ambiguous inputs. The inherent noise introduced through this process also helps to increase model robustness against small perturbations in input space, thus making it more adaptable and robust in practical applications.
@@
While mosaic and mixup augmentation techniques have made significant strides in enhancing neural network robustness, they also present notable limitations. Mosaic can sometimes result in overly complex images that confuse the model rather than help it learn, especially when objects from different images do not naturally coexist. Similarly, mixup might lead to ambiguous labels if the images combined are too distinct, potentially misleading the model during training. Furthermore, methods like copy-paste (Ghiasi et al., 2021) and x-paste (Zhao et al., 2023), which rely on precise segmentation masks and may not be as useful with just bounding box annotations, often face challenges in integrating objects seamlessly into new scenes, leading to unnatural edges or misplaced objects. In contrast, outpainting using diffusion models emerges as a more promising avenue. These models can generate highly realistic and contextually varied synthetic data by extending the canvas of existing images, creating entirely new scenarios that maintain the integrity and realism of the original content. Additionally, the availability of pretrained prompt-guided diffusion models allows for efficient deployment, significantly speeding up the development cycle. This method not only addresses the limitations of conventional augmentation techniques but also provides a scalable solution to train more robust and adaptable AI systems.
@@ 3.1 Creating Seed Vehicle Images
The initial step involves gathering a diverse collection of images of vehicles across the desired categories. This process is manual and entails selecting images where vehicles are presented in a near-eye-level perspective, ensuring that each vehicle is well-separated from smaller background vehicles, if present. Additionally, it is crucial that vehicles are fully visible within the frame, as partial representations result in information loss. For this analysis, our collection comprises over 11,000 vehicle images from various sources (including Stanford Cars) (Krause et al., 2013; Peng et al., 2019; Addison Howard, 2018; POB, 2023; Hoekstra, 2024a, b; Zatoichi, 2023; Koul, 2024) which are utilized exclusively for academic and non-commercial research purposes.
@@
We select the aforementioned detection model by a consensus voting approach, where each model in a predefined ensemble—FCOS (Tian et al., 2019), RetinaNet (Lin et al., 2017), SSD (Liu et al., 2016), MaskRCNN (He et al., 2017), and FasterRCNN (Ren et al., 2016)—detects the largest bounding box for vehicles in a given image. These detections are then compared pairwise using the Intersection over Union (IoU) metric, where each bounding box gains a vote for every other box it significantly overlaps with (as defined by a threshold of 0.95). The model with the highest votes is selected as the primary detector, with the order indicating backup priorities in case the primary model fails. This method leverages multiple models to increase the reliability of the detection by confirming the presence and location of objects through mutual agreement.